The following is a list of financial statement items from the records of Marshall Corporation as at 31 December 2008. Prepaid Insurance $ 12,750 Prepaid Rent $ 18,000 Interest Receivable 0 Salaries Payable 0 Unearned fees $ 30,000 Interest Income $ 10,000

 .
Answer #1
Adjusting entries:
Ref. Particulars Debit Credit
(a) Insurance expense (Note:1) 9000
Prepaid insurance 9000
(Insurance expired)
(b) Interest receivable 2500
Interest income 2500
(Interest accrued)
(c ) Unearned fees (30000/3) 10000
Service revenue 10000
(Fees earned during the year)
(d) Rent expense (18000/9)*5 10000
Prepaid rent 10000
(Rent expired for Aug to Dec-5 months)
(e ) Salaries expense 7000
Salaries payable 7000
(Salaries owed)
Note:1
Inurance expense originally paid for 24 months=$ 18000
Insurance expense per month=18000/24=$ 750
As on Dec 31,2008, only 5 months remain on the 24 months policy
Hence,insurance expense for 12 months expired in 2008=12*750=$ 9000
Prepaid insurance Prepaid rent
Ref. Debit Ref. Credit Ref. Debit Ref. Credit
Bal. 12750 (a) 9000 Bal. 18000 (d) 10000
12750 9000 18000 10000
Bal. 3750 Bal. 8000
Interest receivable Salaries payable
Ref. Debit Ref. Credit Ref. Debit Ref. Credit
(b) 2500 (e ) 7000
Unearned fees Interest income
Ref. Debit Ref. Credit Ref. Debit Ref. Credit
(c ) 10000 Bal. 30000 Bal. 10000
10000 30000 (b) 2500
Bal. 20000 0 12500
Bal. 12500
Service revenue Insurance expense
Ref. Debit Ref. Credit Ref. Debit Ref. Credit
(c ) 10000 (a) 9000
Rent expense Salaries expense
Ref. Debit Ref. Credit Ref. Debit Ref. Credit
(d) 10000 (e ) 7000
Balances:
$
Prepaid insurance 3750
Prepaid rent 8000
Interest receivable 2500
Salaries payable 7000
Unearned fees 20000
Interest income 12500

\